write the equation of the line that contains the indicated pair of points. Express the final equation in standard form. (2,5),(8,12)
Answer by stanbon(75887)

write the equation of the line that contains the indicated pair of points. Express the final equation in standard form. (2,5),(8,12)
-----
slope = (12-5)/(8-2) = 7/6
------------
Form: y = mx + b
Substitite and solve for "b":
5 = (7/6)2+b
b = 5-(7/3)
b = (15/3)-(7/3)
b = 8/3
=========
Equation:
y = (7/6)x+(8/3)
---------
Cheers,
Stan H.